# Human-In-the-Loop Software Development Agents

### v1

@@ Abstract
Recently, Large Language Models (LLMs)-based multi-agent paradigms for software engineering are introduced to automatically resolve software development tasks (e.g., from a given issue to source code). However, existing work is evaluated based on historical benchmark datasets, does not consider human feedback at each stage of the automated software development process, and has not been deployed in practice. In this paper, we introduce a Human-in-the-loop LLM-based Agents framework (HULA) for software development that allows software engineers to refine and guide LLMs when generating coding plans and source code for a given task. We design, implement, and deploy the HULA framework into Atlassian JIRA for internal uses. Through a multi-stage evaluation of the HULA framework, Atlassian software engineers perceive that HULA can minimize the overall development time and effort, especially in initiating a coding plan and writing code for straightforward tasks. On the other hand, challenges around code quality are raised to be solved in some cases. We draw lessons learned and discuss opportunities for future work, which will pave the way for the advancement of LLM-based agents in software development.
@@ I Introduction
Multi-agent paradigm, powered by Large Language Models (LLMs), has demonstrate remarkable results in many research areas [1], including software engineering [2, 3, 4]. Within the software engineering context, LLM-based multi-agent is defined as an autonomous agent (e.g., an LLM model) receiving inputs from surrounding environment (e.g., compilers, linter tools) and performing actions (e.g., writing code, fixing bugs) to achieve a specific objective (e.g., resolving a pull request, passing test cases). The agent will be assigned with a role (e.g., software engineer, software tester) as part of the software development process with access to external resources (e.g., compilers, linters). With this LLM-based multi-agent paradigm, they are able to solve complex software development tasks for a given input and a goal.
Recently, various LLM-based software development agent frameworks are proposed to address complex software development tasks [5, 6, 7, 8, 9, 10, 11]. However, existing LLM-based software development work is (1) evaluated based on historical benchmark datasets of open-source software projects, (2) does not consider human feedback at each stage of the automated software development process, and (3) has not been deployed in practice. These challenges raise important questions of whether the effectiveness of LLM-based software development agent can be generalised beyond the context of open-source software projects, how acceptable are the plans and code generated by LLM-based software development agent at each stage of the software development process, and how do practitioners perceive the benefits and challenges of using LLM-based software development agent.
@@ 1. RQ1)
    For SWE-bench, AI Planner Agent can correctly identify files that need to be changed, achieving an average recall of 86% per issue, while AI Coding Agent can correctly generate code that is similar to the human-written code (i.e., ground-truth) with an average similarity score of 45%. On the other hand, when apply Agents to the Atlassian internal dataset, AI Planner Agent achieves an average recall of 30% per issue, while AI Coding Agent achieves an average similarity score of 30%. We observed that our JIRA issues tend to be shorter than the open-source software projects’.
@@ 2. RQ2)
    Based on the online evaluation of HULA, plans are successfully generated for 527 out of the 663 real-world JIRA issues. Of these, practitioners approved the generated plans for 433 out of 527 plan generated issues, achieving a plan approval rate of 82%. During the coding stage, pull requests were created for 95 out of 376 issues by HULA, achieving a raised PR rate of 25%. Of these, 56 pull requests were successfully merged, leading to a merged PR rate of 59%. These results highlights the success of the HULA framework where human feedback is incorporated in practice.
@@ 3. RQ3)
    The majority of participants agreed that the coding plan was accurate and the generated code was easy to read and modify, which helped reduce their initial development time and effort. However, human involvement is still needed to ensure the code completely addresses the task. In addition, participants acknowledged that the use of HULA promotes good documentation practice to improve the quality of generated plans and codes.
@@
Paper Organization. The paper is organized as follows. Section II describes the software developement workflow at atlassian. Section III discusses related works and limitations. Section IV presents HULA framework. Section V presents the multi-stage evaluation framework. Section VI presents our results. Section VII discusses the lessons learned and future directions. Section VIII discloses the threats to validity. Finally, Section IX draws the conclusion.
@@ II Software Development Workflow at Atlassian
Atlassian is an enterprise software company that develops products for software development, project management, team collaboration, and many more. Atlassian offers tools such as JIRA Software (for task management), Confluence (for collaboration and knowledge management), and Bitbucket (for source code management) to enhance software teams’ productivity, streamline workflows, and facilitate Agile methodologies across various industries. These products serve over 300,000 customers globally, spanning various industries such as technology, finance, healthcare, transportation, and government.
@@
Since these processes are generally done manually, the teams are prone to productivity challenges, e.g., an overwhelming backlog of JIRA issues and context switching between tools. These challenges can then result in delays and inefficiencies in software delivery, prompting Atlassian teams to seek automated solutions that can streamline workflows and improve their overall productivity.
@@ III Related Works and Limitations
Automated software development tools and techniques, powered by deep learning (DL) or large language models (LLMs), are proposed to assist software engineers in various software development tasks (e.g., bug localization [12], code completion [13], code generation [10, 8], testing case generation [14, 15], code review automation [16, 17], vulnerability detection and repair [18, 19]). However, these automated software development tools and techniques are mainly designed to tackle individual SE tasks, leading to suboptimal performance when apply to real-world software development tasks.
Recently, LLM-based multi-agent paradigm for software engineering has emerged [2]. The multi-agent paradigm for software engineering involves multiple autonomous agents, each with specific roles (e.g., software engineer, testers) and goals (e.g., write code, write test cases), working together to solve complex problems. Various LLM-based software development agent frameworks are proposed to resolve complex software development tasks [5, 6, 7, 8, 9, 10, 11]. For example, Yang et al. [5] proposed SWE-agent, an LLM-based agent which build on top of Linux shell, enabling the agent to effectively search, navigate, edit, and execute code commands. Zhang et al. [6] proposed AutoCodeRover, an LLM-based agent with code search on AST (i.e., Abstract Syntax Tree) and spectrum-based fault localization using test suite. Ma et al. [11] proposed RepoUnderstander, an LLM-based agent with Monte Carlo tree search based exploration strategy on repository knowledge graph. Tao et al. [8] proposed Magis, a multi-agent framework with four customized roles: manager, repository custodian, developer, and quality assurance engineer. Chen et al. [7] proposed CodeR, a multi-agent framework with pre-defined task graphs. Arora et al. [10] proposed Masai, a modular framework divided the problem into multiple sub-problems for LLM-based agent.
Limitations. However, existing LLM-based software development work is (1) evaluated based on historical benchmark datasets, limiting the understanding of its effectiveness in the context of resolving enterprise software development tasks; (2) does not consider human feedback at each stage of the automated software development process, limiting the understanding of the practitioners’ acceptability of the plans and code generated by LLM-based software development agent at each stage to resolve enterprise software development tasks; and (3) has not been deployed in practice, limiting the understanding of practitioners’ perceptions on the benefits and challenges of using LLM-based software development agent.
@@ IV HULA: Human-in-the-loop LLM-based Agents Framework for Software Development
At Atlassian, we highly value human intelligence (i.e., the expertise of practitioners) and human authority (i.e., the human abilities to rule AI agents). Therefore, rather than aiming to fully automate software development tasks, we designed an LLM-based software development agent to collaborate with practitioners, functioning as an assistant to help resolve software development tasks, promoting the Human-AI synergy [20, 21]. We envision that in the future of software development, practitioners and LLM-based software development agents should work collaboratively to complete software development tasks. Thus, we design our HULA framework to incorporate human feedback at each step of the software development tasks. This allow practitioners to provide feedback and guidance as needed, ensuring the software development process remains collaborative rather than entirely reliant on fully autonomous agents.
@@ IV-A Framework
  • AI Planner Agent ($\mathcal{P}$) is an LLM model receiving inputs (i.e., a JIRA issue and the source code repository) to identify the more relevant files that are associated with the given JIRA issue (i.e., file localization). Then, the LLM model uses the list of relevant files and the JIRA issue to generate a coding plan (e.g., “add tests for the add_subfigure method”).
@@
  • Human Agent ($\mathcal{H}$) is a software engineer assigned for the given task. The human agent will provide feedback and work cooperatively with AI agents to achieve the common goal (i.e., resolve a JIRA issue).
With the Human-AI synergy in mind, we leverage a human-AI agent coordination paradigm, exploiting human intelligence in each stage to enhance the effectiveness of the multi-agent system for software development and facilitate better alignment with human preference. Different from the other multi-agent systems that leverage competitive, mixed, or hierarchical coordination style without human involvement, our framework exploits the cooperative communication style with the Decentralized Planning Decentralized Execution (DPDE) paradigm to ensure that each AI agent is independently responsible for its own objective and capabilities, while being able to work cooperatively with human agents to achieve the common goal. Although we exploit DPDE paradigm (i.e., no communication among agents), all Agents still have access to the shared memory of information (i.e, the JIRA issue and the source code repository). The merit of this approach lies in minimal communication overhead, reducing the computational resources on the central LLM, enhancing the Agents’ adaptability, as each agent can quickly adjust its behavior based on local information and human feedback. This makes our framework better suited for dynamic and unpredictable environments, while promoting human-AI synergy.
@@ V The Multi-Stage Evaluation of HULA in Real-World Deployment
in Real-World Deployment
In this section, we introduce the multi-stage evaluation framework for evaluating our HULA in real-world deployment. Our evaluation aims at answering the following three research questions.
@@
To answer these research questions, our evaluation consists of three stages, i.e, offline evaluation, online evaluation, and practitioner survey. The main purpose of offline evaluation is to validate the performance of HULA prior to the actual deployment. Since the offline evaluation did not incorporate human feedback and cannot capture the real-world usage, we conduct an online evaluation to investigate how practitioners use HULA in practice. However, the online evaluation still lacks the user perception. To address this, we run a practitioner survey to investigate the performance of HULA from user perspectives. Below, we present the research methodology for each stage of the evaluation (see Figure 3).
@@ V-A Stage 1: Offline Evaluation
The offline evaluation based on benchmark datasets allows us to preliminary assess the performance of HULA to ensure that it achieves an acceptable performance before a deployment. Note that we did not include human feedback for this offline experiment to fully automate the evaluation process.
Benchmark Datasets: To conduct an offline evaluation, we need high-quality benchmark datasets that provide sufficient information of the work to be done and human-written code solutions for diverse programming languages. Thus, we use the SWE-bench dataset which provides high-quality datasets for Python programming language. Then, we collected development issues at Atlassian to evaluate the effectiveness of HULA in an industrial setting that also cover diverse context and programming languages. Table I provides an overview of our benchmark datasets.
@@
  • Internal dataset is a collection of completed development tasks (i.e., JIRA issues) at Atlassian. We randomly sampled issues from JIRA Software with the corresponding pull requests (PRs) from BitBucket. In total, our internal dataset contains 369 JIRA issues from 94 software repositories. The issues’ context ranges from front-end developments, back-end developments, data analytics, security patches, product migrations, configuration updates, bug fixes, and more. The code solutions in PRs involve various file types with more than 10 programming languages including TypeScript, Java, Kotlin, Go, Python.
@@
  • %Issues for Success Generation measures the percentage of issues that HULA can successfully generate plans and code, completing all of the steps in the workflow (see Figure 1) (i.e., success generation). Otherwise, if HULA fails at any step in the workflow and unable to produce the final code changes, we consider it as unsuccessful.
@@
  • %Issues of High Code Similarity measures the percentage of issues that the AI generated code is highly similar to the ground-truth human-written code. Code similarity is used as a proxy to determine if the code being generated is highly similar to the ground-truth human-written code, due to the absence of the test cases in our internal dataset. Typically, human involvement is typically required to determine if the code is similar or not, yet it is very expensive in practice. Recently, CodeBERTScore [25] or BERTScore [26] are widely-used in software engineering to evaluate code similarity. However, these techniques are designed for BERT-like architectures, limiting its input length to only 512 tokens, which does not cover the typical size of the human-written code (i.e., a median of 1,275 tokens for our internal dataset). Following the LLM-as-a-Judge principle[27], we exploit GPT-4 as a judge to measure the similarity between the source code and the ground-truth, given the corresponding JIRA issue as a context. After iterations of prompt refinement, we develop a prompt instruction containing $<$prompt, issue, generated code, ground-truth code$>$ to generate the similarity score ranging from 1 (no similarity), 2 (weak similarity), 3 (high similarity), to 4 (highly similarity). Any issue that achieves the code similarity score of 3 to 4 is considered as high code similarity.
@@ V-B Stage 2: Online Evaluation
Internal Deployment: To avoid interruption in day-to-day work of Atlassian practitioners, we opted to deploy this new technology carefully as per the following steps. First, we conducted initial trials with a small group of software engineers from two software development teams to gain initial feedbacks. After ensured positive feedback with the first group, we gradually deploy HULA with a broader group of practitioners. Figure 2 shows an example of HULA workflow that integrated into a JIRA issue.
@@
  • Code Generation Rate measures the percentage of issues that HULA can complete the workflow and generate code changes, compared to the total number of issues that with attempted code generation.

### v2

@@ Abstract
Recently, Large Language Models (LLMs)-based multi-agent paradigms for software engineering are introduced to automatically resolve software development tasks (e.g., from a given issue to source code). However, existing work is evaluated based on historical benchmark datasets, rarely considers human feedback at each stage of the automated software development process, and has not been deployed in practice. In this paper, we introduce a Human-in-the-loop LLM-based Agents framework (HULA) for software development that allows software engineers to refine and guide LLMs when generating coding plans and source code for a given task. We design, implement, and deploy the HULA framework into Atlassian JIRA for internal uses. Through a multi-stage evaluation of the HULA framework, Atlassian software engineers perceive that HULA can minimize the overall development time and effort, especially in initiating a coding plan and writing code for straightforward tasks. On the other hand, challenges around code quality remain a concern in some cases. We draw lessons learned and discuss opportunities for future work, which will pave the way for the advancement of LLM-based agents in software development.
@@ I Introduction
Multi-agent paradigm, powered by Large Language Models (LLMs), has demonstrated remarkable results in many research areas [1], including software engineering [2, 3, 4]. Within the software engineering context, an LLM-based multi-agent is defined as an autonomous agent (e.g., an LLM model) receiving inputs from the surrounding environment (e.g., compilers, linter tools) and performing actions (e.g., writing code, fixing bugs) to achieve a specific objective (e.g., resolving a pull request, passing test cases). The agent will be assigned with a role (e.g., software engineer, software tester) as part of the software development process with access to external resources (e.g., compilers, linters). With this LLM-based multi-agent paradigm, they are able to solve complex software development tasks for a given input and a goal.
Recently, various LLM-based software development agent frameworks are proposed to address complex software development tasks [5, 6, 7, 8, 9, 10, 11]. However, existing LLM-based software development work is (1) evaluated based on historical benchmark datasets of open-source software projects, (2) does not consider human feedback at each stage of the automated software development process, and (3) has not been deployed in practice. These challenges raise important questions of whether the effectiveness of LLM-based software development agents can be generalised beyond the context of open-source software projects, how acceptable the plans and code generated by LLM-based software development agents at each stage of the software development process, and how practitioners perceive the benefits and challenges of using LLM-based software development agent.
@@ 1. RQ1)
    For SWE-bench, AI Planner Agent can correctly identify files that need to be changed, achieving an average recall of 86% per issue, while AI Coding Agent can correctly generate code that is similar to the human-written code (i.e., ground-truth) with an average similarity score of 45%. On the other hand, when applying Agents to the Atlassian internal dataset, AI Planner Agent achieves an average recall of 30% per issue, while AI Coding Agent achieves an average similarity score of 30%. We observed that our JIRA issues tend to be shorter than the open-source software projects.
@@ 2. RQ2)
    Based on the online evaluation of HULA, plans are successfully generated for 527 out of the 663 real-world JIRA issues. Of these, practitioners approved the generated plans for 433 out of 527 plan-generated issues, achieving a plan approval rate of 82%. During the coding stage, pull requests (PRs) were created for 95 out of 376 code-generated issues, achieving a raised PR rate of 25%. Of these, 56 PRs were successfully merged, leading to a merged PR rate of 59%. These results highlight the success of the HULA framework where human feedback is incorporated into practice.
@@ 3. RQ3)
    41% of participants agreed that the generated plan by HULA aligned accurately with the issue description. While HULA still requires human involvement to ensure the generated code completely addresses the task, 61% of the participants agreed that the generated code was easy to read and modify, which helped reduce their initial development time and effort. Additionally, participants acknowledged that HULA promotes good documentation practice, though this requires more effort to provide detailed issue descriptions.
@@
Paper Organization. The paper is organized as follows. Section II describes the software development workflow at Atlassian. Section III discusses related works and limitations. Section IV presents HULA framework. Section V presents the multi-stage evaluation framework. Section VI presents our results. Section VII discusses the lessons learned and future directions. Section VIII discloses the threats to validity. Finally, Section IX draws the conclusion.
@@ II Software Development Workflow at Atlassian
Atlassian is an enterprise software company that develops products for software development, project management, team collaboration, and many more. Atlassian offers tools such as JIRA Software (for task management), Confluence (for collaboration and knowledge management), and Bitbucket (for source code management) to enhance software teams’ productivity, streamline workflows, and facilitate Agile methodologies across various industries. These products serve over 300,000 customers globally, spanning various industries such as technology, finance, healthcare, and government.
@@
In 2024, Atlassian had over 12,000 engineers submitting over 950,000 pull requests. Since these processes are generally done manually, the teams are prone to productivity challenges, e.g., an overwhelming backlog of JIRA issues and context switching between tools. These challenges can result in delays and inefficiencies in software delivery, prompting Atlassian teams to seek automated solutions that can streamline workflows and improve their overall productivity.
@@ III Related Works and Limitations
Automated software development tools and techniques, powered by deep learning (DL) or large language models (LLMs), are proposed to assist software engineers in various software development tasks (e.g., bug localization [12, 13, 14], code completion [15, 16], code generation [10, 8], test case generation [17, 18], code review automation [19, 20, 21, 22], vulnerability detection and repair [23, 24, 25, 26]). However, these automated software development tools and techniques are mainly designed to tackle individual SE tasks, leading to suboptimal performance when applied to real-world software development tasks.
Recently, an LLM-based multi-agent paradigm for software engineering has emerged [2]. The multi-agent paradigm for software engineering involves multiple autonomous agents, each with specific roles (e.g., software engineer, testers) and goals (e.g., writing code, writing test cases), working together to solve complex problems. Various LLM-based software development agent frameworks are proposed to resolve complex software development tasks [5, 6, 7, 8, 9, 10, 11]. For example, Yang et al. [5] proposed SWE-agent, an LLM-based agent built on top of a Linux shell, enabling the agent to effectively search, navigate, edit, and execute code commands. Zhang et al. [6] proposed AutoCodeRover, an LLM-based agent with code search on AST (i.e., Abstract Syntax Tree) and spectrum-based fault localization using a test suite. Ma et al. [11] proposed RepoUnderstander, an LLM-based agent with a Monte Carlo tree search-based exploration strategy on the repository knowledge graph. Tao et al. [8] proposed Magis, a multi-agent framework with four customized roles: manager, repository custodian, developer, and quality assurance engineer. Chen et al. [7] proposed CodeR, a multi-agent framework with pre-defined task graphs. Arora et al. [10] proposed Masai, a modular framework that divided the problem into multiple sub-problems for LLM-based agents.
Limitations. However, existing LLM-based software development work is (1) evaluated based on historical benchmark datasets, limiting the understanding of its effectiveness in the context of resolving enterprise software development tasks; (2) rarely considering human feedback at each stage of the automated software development process, limiting the understanding of the practitioners’ acceptability of the plans and code generated by LLM-based software development agent at each stage to resolve enterprise software development tasks; and (3) has not been deployed in practice, limiting the understanding of practitioners’ perceptions on the benefits and challenges of using LLM-based software development agent.
@@ IV HULA: Human-in-the-loop LLM-based Agents Framework for Software Development
At Atlassian, human intelligence (i.e., the expertise of practitioners) and human authority (i.e., the human ability to rule AI agents) are highly valued. Therefore, rather than aiming to fully automate software development tasks, we designed an LLM-based software development agent to collaborate with practitioners, functioning as an assistant to help resolve software development tasks, promoting the Human-AI synergy [27, 28]. We envision that in the future of software development, practitioners and LLM-based software development agents should work collaboratively to complete software development tasks. Thus, we design our HULA framework to incorporate human feedback at each step of the software development tasks. This allows practitioners to provide feedback and guidance as needed, ensuring the software development process remains collaborative rather than entirely reliant on fully autonomous agents.
@@ IV-A Framework
  • AI Planner Agent ($\mathcal{P}$) is an LLM model receiving inputs (i.e., a JIRA issue and the source code repository) to identify the more relevant files that are associated with the given JIRA issue (i.e., file localization  [13, 14, 29]). Then, the LLM model uses the list of relevant files and the JIRA issue to generate a coding plan (e.g., “add tests for the add_subfigure method”).
@@
  • Human Agent ($\mathcal{H}$) is a software engineer assigned to the given task. The human agent will provide feedback and work cooperatively with AI agents, particularly, reviewing and editing the plan and the code, and raising PRs, to achieve the common goal (i.e., resolve a JIRA issue).
With the Human-AI synergy in mind, we leverage a human-AI agent coordination paradigm, exploiting human intelligence in each stage to enhance the effectiveness of the multi-agent system for software development and facilitate better alignment with human preference. Different from the other multi-agent systems that leverage competitive, mixed, or hierarchical coordination styles without human involvement, our framework exploits the cooperative communication style with the Decentralized Planning Decentralized Execution (DPDE) paradigm to ensure that each AI agent is independently responsible for its own objective and capabilities while being able to work cooperatively with human agents to achieve the common goal. Although we exploit the DPDE paradigm (i.e., no communication among agents), all Agents still have access to the shared memory of information (i.e., the JIRA issue and the source code repository). The merit of this approach lies in minimal communication overhead, reducing the computational resources on the central LLM, and enhancing the Agents’ adaptability, as each agent can quickly adjust its behaviour based on local information and human feedback. This makes our framework better suited for dynamic and unpredictable environments while promoting human-AI synergy.
@@ V The Multi-Stage Evaluation of HULA in Real-World Deployment
In this section, we introduce the multi-stage evaluation framework for evaluating our HULA in real-world deployment. Our evaluation aims to answer the following three research questions.
@@
To answer these research questions, our evaluation consists of three stages, i.e., offline evaluation, online evaluation, and practitioner survey. The main purpose of offline evaluation is to validate the performance of HULA prior to the actual deployment. Since the offline evaluation did not incorporate human feedback and cannot capture the real-world usage, we conducted an online evaluation to investigate how practitioners use HULA in practice. However, the online evaluation still lacks the user perception. To address this, we run a practitioner survey to investigate the performance of HULA from user perspectives. Below, we present the research methodology for each stage of the evaluation (see Figure 3).
@@ V-A Stage 1: Offline Evaluation
The offline evaluation based on benchmark datasets allows us to preliminary assess the performance of HULA to ensure that it achieves an acceptable performance before deployment. Note that we did not include human feedback for this offline experiment to fully automate the evaluation process.
Benchmark Datasets: To conduct an offline evaluation, we need high-quality benchmark datasets that provide sufficient information of the work to be done and human-written code solutions for diverse programming languages. Thus, we use the SWE-bench dataset which provides high-quality datasets for Python programming language. Then, we collected development issues at Atlassian to evaluate the effectiveness of HULA in an industrial setting that also covers diverse contexts and programming languages. Table I provides an overview statistics of our benchmark datasets.
@@
  • Internal dataset is a collection of completed development tasks (i.e., JIRA issues) at Atlassian. We randomly sampled issues from JIRA Software with the corresponding pull requests (PRs) from BitBucket. In total, our internal dataset contains 369 JIRA issues from 94 software repositories. The issues’ context ranges from front-end developments, back-end developments, data analytics, security patches, product migrations, configuration updates, bug fixes, and more. The code solutions in PRs involve various file types with more than 10 programming languages including TypeScript, Java, Kotlin, and Python.
@@
  • %Issues for Success Generation measures the percentage of issues that HULA can successfully generate plans and code, completing all of the steps in the workflow (see Figure 1) (i.e., success generation). Otherwise, if HULA fails at any step in the workflow and is unable to produce the final code changes, we consider it as unsuccessful.
@@
  • %Issues of High Code Similarity measures the percentage of issues that the AI-generated code is highly similar to the ground-truth human-written code. Code similarity is used as a proxy to determine if the code being generated is highly similar to the ground-truth human-written code, due to the absence of the test cases in our internal dataset. Typically, human involvement is typically required to determine if the code is similar or not, yet it is very expensive in practice. Recently, CodeBERTScore [33] or BERTScore [34] are widely used in software engineering to evaluate code similarity. However, these techniques are designed for BERT-like architectures, limiting its input length to only 512 tokens, which does not cover the typical size of the human-written code (i.e., a median of 1,275 tokens for our internal dataset). Following the LLM-as-a-Judge principle[35], we exploit GPT-4 as a judge to measure the similarity between the source code and the ground-truth, given the corresponding JIRA issue as a context. After iterations of prompt refinement, we develop a prompt instruction containing $<$prompt, issue, generated code, ground-truth code$>$ to generate the similarity score ranging from 1 (no similarity), 2 (weak similarity), 3 (high similarity), to 4 (highly similarity). Any issue that achieves a code similarity score of 3 to 4 is considered high code similarity.
@@ V-B Stage 2: Online Evaluation
Internal Deployment: To avoid interruption in the day-to-day work of Atlassian practitioners, we opted to deploy this new technology carefully as per the following steps. First, we conducted initial trials with a small group of software engineers from two software development teams to gain initial feedback. After ensuring positive feedback from the first group, we gradually deploy HULA with a broader group of practitioners. Figure 2 shows an example of HULA workflow that integrated into a JIRA issue.
@@
  • Code Generation Rate measures the percentage of issues that HULA can complete the workflow and generate code changes, compared to the total number of issues with attempted code generation.